OpenStudy (ajoy31):

how do I factor x^3+3x^2-6x-8?

OpenStudy (anonymous):

(x+1)(x+4)(x-2) = 0

OpenStudy (anonymous):

First I would re-arrange and look for patterns. Noting that x^3-8 looks pretty perfect cubes to me. using (a^3-b^3)=(a-b)(a^2+ab+b^2) we may get somewhere. re-arrange x^3-8+3x^2-6x group first two and second two terms and factor (x-2)(x^2+2x+4)+3x(x-2) factor out x-2 (x-2)(x^2+2x+4+3x) combine like terms (x-2)(x^2+5x+4) then factor the poynomial (x-2)(x+1)(x+4)